OpenStudy (ny,ny):

if angle MOL and angle LON form a linear pair, what is m angle MOL if m angle MOL is 3x+2 and m angle LON = 10x-4. is line LO an altitude?

OpenStudy (anonymous):

linear pair implis sum is 180 degrees (3x+2)+(10x-4)=180 13x-2=180 x=182/13= 14

OpenStudy (ny,ny):



OpenStudy (anonymous):

m angle MON= 180 = m angle MOL+ m angle LON= 3x+2 + 10x-4

OpenStudy (anonymous):

3x+2=3*14+2=42+2=44 not its not altitude as its not right angle

OpenStudy (ny,ny):

okay thank you